OpenStudy (anonymous):

what is one ordered pair that is a solution of 8x-5y=-2?

OpenStudy (anonymous):

what are the choices?

OpenStudy (anonymous):

There are none.. :c It just says identify one ordered pair.

jimthompson5910 (jim_thompson5910):

For each ordered pair, you just plug it into the equation and evaluate/simplify. For instance, if you are given the ordered pair (0,3), you plug in x = 0 and y = 3 to get 8x-5y=-2 8*0-5*3=-2 0 - 15 = -2 -15 = -2 ... this equation is FALSE Since the equation above is FALSE, this means that (0,3) is NOT a solution to 8x-5y=-2

jimthompson5910 (jim_thompson5910):

If you don't have any ordered pairs, then pick any number you want for x and plug it in for x. Then solve for y to get the corresponding value

OpenStudy (anonymous):

Okay. Thank you.

jimthompson5910 (jim_thompson5910):

Sure thing